# On the minimax optimality of Flow Matching through the connection to kernel density estimation

### v1

@@ 3.2 Flow Matching with parameterized vector field
3.3 Density estimation on a linear subspace
The rate in Theorem 3.6 depends on the dimension $d$ and therefore exhibits the classical curse of dimensionality. In this section, we analyze a case, in which the curse of dimensionality can be circumvented and the rate depends on some intrinsic dimension $d^{\prime}<d$. Assume that the support of $\mathbb{P}^{*}$ is a subset of an $d^{\prime}$ dimensional linear subspace of $\mathbb{R}^{d}$. Let the linear subspace $V$ be defined by
$V\coloneqq\{x\in\mathbb{R}^{d}\colon\exists y\in\mathbb{R}^{d^{\prime}}\text{ s.t. }x=Ay\},$
where $A$ is a $\mathbb{R}^{d\times d^{\prime}}$ matrix with normalized orthogonal columns. Let $p^{*}_{d^{\prime}}$ be the density with respect to the Lebesgue measure on $\mathbb{R}^{d^{\prime}}$, such that for $Y\sim p^{*}_{d^{\prime}}$ we have that $AY\sim p^{*}$, where $p^{*}$ is the density of $\mathbb{P}^{*}$ with respect to the Lebesgue measure restricted on $V$. This setting has been studied in context of diffusions by Chen et al. [2023b] and Oko et al. [2023].
In this case we can improve the rate of convergence, such that the corresponding rate depends only on the intrinsic dimension $d^{\prime}$ and not on the ambient dimension $d.$ We focus on the case of the Gaussian kernel.
@@ 3.3 Density estimation on a linear subspace
Assume $d^{\prime}\geq 2,p^{*}\in B_{1,\infty}^{\alpha},\alpha\in(0,1]$ with $\|p^{*}\|_{B_{1,\infty}^{\alpha}}\leq M$, $\operatorname{supp}(p^{*})$ bounded and $\mathbb{U}=\mathcal{N}(0,I_{d})$. Then there are $C_{1}$ and $C_{2}$ such that for $v^{n}$ from (8)
$\mathbb{E}\big{[}\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi^{n}_{1}(Z)})\big{]}\leq C_{1}\sigma_{\min}^{1+\alpha}+\frac{C_{2}}{\sqrt{\sigma^{d^{\prime}-2}}}\frac{\log n}{\sqrt{n}}.$
For $\sigma_{\operatorname{min}}\sim(n/\log^{2}n)^{-\frac{1}{2\alpha+d^{\prime}}}$ we obtain $\mathbb{E}\big{[}\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi^{n}_{1}(Z)})\big{]}=\mathcal{O}((n/\log^{2}n)^{-\frac{1+\alpha}{2\alpha+d^{\prime}}}).$
The above result is in line with the similar results for kernel density estimators for distributions on manifolds Berenfeld and Hoffmann [2021], Divol [2022]. Note however that in these articles, the kernel is normalized with respect to the lower dimension, i.e. they consider $h^{-d^{\prime}}K(\cdot/h)$ for bandwidth $h>0$, while our generative setting enforces us to prove Theorem 3.8 for kernels which are normalized with respect to the ambient space dimension, i.e. $h^{-d}K(\cdot/h)$ (and $h=\sigma_{\min}$ in our notation).
Analogous to Theorem 3.6, we obtain a rate of convergence in case ReLU networks are used for the vector field.
@@
Grant Assumption 3.5 and assume $\operatorname{supp}(p^{*})\subset[-1,1]^{d}$ and $p^{*}\in B^{\alpha}_{1,\infty}$ with $\|p^{*}\|_{B^{\alpha}_{1,\infty}}\leq M<\infty,\alpha\in(0,1]$. Set $\sigma_{\min}=n^{-\frac{1}{2\alpha+d^{\prime}}}.$ Then there are sequences $L_{n},M_{n},\Gamma_{n}\in\mathbb{N}$ such that for $n$ big enough
$\mathbb{E}\big{[}\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\hat{\psi}_{1}})\big{]}\lesssim n^{-\frac{1+\alpha}{2\alpha+d^{\prime}}}\log^{2}(n),$
where $\hat{\psi}$ is given by (10) with $\mathcal{M}=\mathsf{NN}(L_{n},M_{n},\Gamma_{n})$.
Theorem 3.8 and Theorem 3.9 demonstrate that Flow Matching can benefit from lower intrinsic dimensions. Note that the method automatically adapts to the linear subspace, only the dimension $d^{\prime}$ is required for the choice of $\sigma_{\min}$. However, the smaller error bound comes at the cost of even larger networks to ensure that the approximation error of the network remains negligible compared to the faster rate of convergence in the dominating first term in decomposition (15). Therefore, the above results can only serve as first step to an analysis in a dimension reduction setting. A natural next step would be a generalization from linear subspaces to $d^{\prime}$-dimensional sub-manifolds. Moreover, it would be desirable to achieve the same rate with a smaller network, ideally even benefiting from the smaller dimension.
@@ 4 Proofs
Before we begin with the proofs, let us introduce some additional notation: The spectral norm of $A$ is denoted by $\|A\|_{2}.$ We define the ceiling of $x\in\mathbb{R}$ as $\lceil x\rceil:=\min\{k\in\mathbb{Z}|k\geq x\}.$ For $\alpha\in(0,1]$, we define the Hölder space $\mathcal{H}^{\alpha}$ by
$\mathcal{H}^{\alpha}(\Omega)\coloneqq\big{\{}f\colon\Omega\rightarrow\mathbb{R}^{d}\colon\|f\|_{\mathcal{H}^{\alpha}(\Omega)}<\infty\big{\}},\quad\text{ where }\quad\|f\|_{\mathcal{H}^{\alpha}(\Omega)}\coloneqq\sup_{x,y\in\Omega}\frac{|f(x)-f(y)|}{|x-y|^{\alpha}}.$
@@
4.1 Proofs of Section 2
Proof of Lemma 2.1
Proof.
@@ 4.1 Proofs of Section 2
$\mathbb{E}_{X_{t}\sim p_{t}}[\langle\hat{v}_{t}(X_{t}),v^{n}_{t}(X_{t})\rangle]$ $=\int\langle\hat{v}_{t}(x),v^{n}_{t}(x)\rangle p_{t}(x)\;\mathrm{d}x$
@@
4.2 Proofs and extended results of Section 3
Proof of Theorem 3.1
Proof.
As $\mathsf{W}_{1}$ and $\mathsf{W}_{2}$ satisfy the triangle inequality (see Villani [2008, p. 94]),

### v2

@@ 3.2 Flow Matching with parameterized vector field
3.3 Density estimation on a manifold
The rate in Theorem˜3.7 depends on the dimension $d$ and therefore exhibits the classical curse of dimensionality. In this subsection, we analyze a case, in which the curse of dimensionality can be circumvented and the rate depends on some intrinsic dimension $d^{\prime}<d$.
To this end, we consider a single-chart $d^{\prime}$-dimensional manifold $\mathcal{M}$, i.e. there exists a smooth, invertible map $g\colon\mathbb{R}^{d^{\prime}}\rightarrow\mathbb{R}^{d}$ such that
$\mathcal{M}=g(\mathbb{R}^{d^{\prime}}).$
This setting is commonly assumed in generative modeling, specifically in GANs, see for example Schreuder et al. (2021). For our analysis, we impose the following assumption on the smoothness of $g$.
Assumption 3.9.
Let $g\in C^{2}$ and $\|g\|_{\infty}\leq M,\|Dg(y)\|\leq M_{1},\|D(Dg(y))\|\leq M_{2}$. The Jacobian $Dg$ is such that for every $y\in\mathbb{R}^{d^{\prime}}$, $0<m<\lambda_{\min}(Dg(y))<\lambda_{\max}(Dg(y))<M_{3}$, where $\lambda_{\min}$ is the smallest and $\lambda_{\max}$ is the largest eigenvalue and $m,M_{3}>0$ are constants. The inverse $g^{-1}$ of $g$ is twice continuously differentiable.
We further assume that $\mathcal{M}$ has a positive reach $\tau>0$. The notion of the reach of a manifold is due to Federer (1959). First, we define the distance function to $\mathcal{M}$ for $x\in\mathbb{R}^{d}$ as $\operatorname{dist}(x,M)=\inf_{z\in\mathcal{M}}|x-z|$. Then, the medial axis of $\mathcal{M}$ is given by
$\operatorname{Med}(\mathcal{M})=\{x\in\mathbb{R}^{d}\mid\exists\ y,z\in\mathcal{M},y\neq z,|y-x|=|z-x|=\operatorname{dist}(x,\mathcal{M})\}.$
The reach of $\mathcal{M}$ is then defined as
$\tau=\inf_{x\in\mathcal{M}}\operatorname{dist}(x,\operatorname{Med}(\mathcal{M})).$
For an illustration of the concept of the reach see Figure˜4 in Section˜5.3. For $r<\tau$ and $r<c_{0}=\frac{2m^{2}\tau}{2M_{2}\tau+3mM_{1}}$, let $\pi\colon\mathbb{R}^{d}\rightarrow\mathcal{M}\cup\{0\}$ be a function such that $\pi\big|_{B(r,\mathcal{M})}$ is the orthogonal projection onto $\mathcal{M}$ and $\pi(B(r,\mathcal{M})^{c})=\{0\}$. In this subsection, we assume that $\mathbb{P}^{*}$ is supported on $\mathcal{M}$ and that there is a $\alpha$-Besov smooth distribution $\mathbb{P}^{*}_{d^{\prime}}$ on $\mathbb{R}^{d^{\prime}}$ such that $g(Y)\sim\mathbb{P}^{*}$ when $Y\sim\mathbb{P}^{*}_{d^{\prime}}$. The next theorem bounds the estimation error of the KDE in a proximity of the manifold.
@@ 3.3 Density estimation on a manifold
Assume $\mathbb{P}^{*}_{d^{\prime}}$ has a Lebesgue density $p^{*}_{d^{\prime}}\in B^{\alpha}_{1,\infty}(\Lambda,[-1,1]^{d^{\prime}})$ for $\alpha\in(0,1],\Lambda>0$. Further assume that the kernel $K$ satisfies the assumptions of Theorem˜3.4. Then
$\mathbb{E}\big[\mathsf{W}_{1}(\mathbb{P}^{*},\pi_{\#}\mathbb{P}^{\psi^{n}_{1}(Z)})\big]\leq C_{1}\sigma_{\min}^{1+\alpha}+\frac{C_{2}}{\sqrt{\sigma_{\min}^{d^{\prime}-2}}}\frac{\log n}{\sqrt{n}}.$
For $\sigma_{\min}\sim(n/\log^{2}(n))^{-\frac{1}{2\alpha+d^{\prime}}}$ we obtain $\mathbb{E}\big[\mathsf{W}_{1}(\mathbb{P}^{*},\pi_{\#}\mathbb{P}^{\psi^{n}_{1}}(Z))\big]=\mathcal{O}\big((n/\log^{2}(n))^{-\frac{1+\alpha}{2\alpha+d^{\prime}}}\big)$.
The proof of Theorem˜3.10 is inspired and uses techniques from Aamari et al. (2019) and Genovese et al. (2014). The proof starts with a decomposition of the projection into tangential vectors and a remainder. The remainder is bounded using a Taylor series and the reach assumption. The rest of the proof follows along the lines of Theorem˜3.4, but is carefully adapted to the manifold and projection setting. While the rigorous arguments are very technical, in a nutshell, the tangential noise is retraced to $\mathbb{R}^{d^{\prime}}$ and the space dependence of this retraction is shown to be small. Just as Theorem˜3.4, the proof strongly relies on the dual form of the Wasserstein $1$ metric, thus extensions to Wasserstein $2$ are not straightforward.
The above result is in line with the similar results for kernel density estimators for distributions on manifolds, c.f. Berenfeld and Hoffmann (2021); Divol (2022). Note however that in these articles, the kernel is normalized with respect to the lower dimension, i.e. they consider $h^{-d^{\prime}}K(\cdot/h)$ for bandwidth $h>0$, while our generative setting enforces us to prove Theorem˜3.10 for kernels which are normalized with respect to the ambient space dimension, i.e. $h^{-d}K(\cdot/h)$ (and $h=\sigma_{\min}$ in our notation). Furthermore, in Divol (2022, Theorem 3.1), the kernel estimator is evaluated on the manifold. Since the Flow Matching estimator results in generated samples and we evaluate the distribution of these samples, we cannot measure the Wasserstein distance along the manifold without the tube-projection, since the density of a distribution restricted to a null-set is not unique. This is different from the density estimation settings, where the density is either constructed directly on the manifold (Divol, 2022) or evaluated pointwise (Berenfeld and Hoffmann, 2021). Section˜4 will give another motivation for this evaluation measure.
@@
Grant Assumption 3.6 and assume the setting of Theorem˜3.10. Set $\sigma_{\min}=n^{-\frac{1}{2\alpha+d^{\prime}}}.$ Then there are sequences $L_{n},S_{n},\Gamma_{n}\in\mathbb{N}$ such that for $n$ big enough
$\mathbb{E}\big[\mathsf{W}_{1}(\mathbb{P}^{*},\pi_{\#}\mathbb{P}^{\hat{\psi}_{1}(Z)})\big]\lesssim n^{-\frac{1+\alpha}{2\alpha+d^{\prime}}}\log(n)^{2},$
where $\hat{\psi}$ is given by (11) with $\mathcal{N}=\mathsf{NN}(L_{n},M_{n},\Gamma_{n})$.
Theorem˜3.11 shows that the Flow Matching estimator can approximate $\mathbb{P}^{*}$ with a faster rate on a small tube around the manifold, exploiting the smoothness of $\mathbb{P}^{*}$ on the manifold. It uses exactly the same argument as Theorem˜3.7. For diffusion models, similar behavior has been observed by Tang and Yang (2024); Azangulov et al. (2024).
4 Numerical illustration
While the empirical success is widely known, the main argument of this work is that in an overparameterized setting the Flow Matching algorithm is at least as good as the kernel density estimator. In a full dimensional setting, we still recover a minimax optimal rate of convergence. Particularly, the Flow Matching algorithm does not mimic the empirical measure, but a smoothed version. This enables leveraging smoothness, even in an overparameterized setting. In this section, we want to illustrate this comparison using a constructed example.
To this end, we consider the following setting: the distribution $\mathbb{P}^{*}$ is the uniform distribution on the graph of the sine curve restricted to [3,3][-3,3]. Thus, $\mathbb{P}^{*}$ lies on a $1$-dimensional manifold in $\mathbb{R}^{2}$. We assume that we have $200$ samples from $\mathbb{P}^{*}$. In this setting, we train a Flow Matching model with three hidden SeLU layers of width $512$ and $\mathbb{U}=\mathcal{N}(0,I_{2})$, for a definition of the SeLU activation function see Klambauer et al. (2017). Our implementation builds up on the code provided by Tong et al. (2024), we use the Adam optimizer (Kingma and Ba, 2014) with the standard parameters $(\operatorname{lr}=0.001$, $\beta_{1}=0.9$ and $\beta_{2}=0.999)$ and Poli et al. (2025) to solve the neural ODE with the solver dopri5 and set $\mathrm{atol}=\mathrm{rtol}=10^{-5}$. In order to analyse the model without avoidable additional smoothing, we do not use minibatches. We evaluated the estimator after 10 000,20 000,30 00010\ 000,20\ 000,30\ 000 and 40 00040\ 000 updates to avoid committing to a fixed number of training iterations. As a comparison, we use a KDE with a Gaussian kernel. For both methods, we generate $512$ samples. To interpret the KDE as a generative model, we draw an index uniformly, sample from $\mathcal{N}(0,I_{2})$ and then add the noise to the sample of the index. An example of this setting and generated samples from the generative models for $\sigma_{\min}\in\{0.5,0.1,0.05,0.01\}$ can be found in Figure˜1. We repeated the following experiments for networks with width of $32$ and $8$, the results can be found in Figure˜5 in Appendix˜A.
(a) FM, $\sigma_{\min}=0.5$
(b) FM, $\sigma_{\min}=0.1$
(c) FM, $\sigma_{\min}=0.05$
(d) FM, $\sigma_{\min}=0.01$
(e) KDE, $\sigma_{\min}=0.5$
(f) KDE, $\sigma_{\min}=0.1$
(g) KDE, $\sigma_{\min}=0.05$
(h) KDE, $\sigma_{\min}=0.01$
Figure 1: Exemplarily result of the Flow Matching (FM) estimator and the KDE. The generated samples are the blue dots, the training sample is represented by black dots on the red sine curve. Top row shows generated samples of the FM estimator after $40000$ runs.
By construction, we know that the KDE does not detect the manifold, but spreads noise in two dimensions. The question is whether the Flow Matching estimator is capable of detecting the manifold, without imposing structural assumptions on the manifold. To measure this, we calculate the mean distance of the generated samples to the manifold. We repeated the same experiment independently drawing new samples $10$ times. The results are gathered in Figure˜2(a). For all choices of $\sigma_{\min}$ except the largest, the KDE achieves the smallest mean distance, which strongly opposes the claim that the Flow Matching estimator detects the manifold better that the KDE.
Naturally, for very small $\sigma_{\min}$, the KDE is very close to the empirical distribution. In proofs, i.e. the proof of Theorem˜3.4 only a proper choice of $\sigma_{\min}$ leads an the optimal rate. The Flow Matching estimator does not behave the same, the model seems to smooth intrinsically. Albeit we ran the model for a very large number of training iterations, there is no mechanism to prevent the convergence towards a local minimum. Therefore this effect could be due to the optimization error. More desirable is an intrinsic smoothing. In order to measure this effect, we calculated the maximal distance between two projections on the sine curves in the $10$ runs. In Figure˜2(b) it is visible that for smaller $\sigma_{\min}$, the Flow Matching estimator approximately covers the manifold better than then KDE, whereas this cannot be seen for larger $\sigma_{\min}$. While for the uniform distribution that is considered here, gaps in the covering are a satisfactory indication for the quality of estimation, this measure fails for other distributions. In hindsight, this motivates the projected measure evaluated in Theorem˜3.10.
(a) Mean distance of the generated samples to the manifold. Note that the shape of the sine curve leads to a smaller expected distance than $\sigma_{\min}$, this effect is smaller for smaller $\sigma_{\min}$.
(b) Largest gap on the manifold between the projections of the generated samples.
Figure 2: Comparison between the Flow Matching estimator and the KDE in terms of distance to the manifold and the largest gap of the projection estimator on the manifold.
Our results indicate that in order to theoretically find situations in which the Flow Matching algorithm outperforms the KDE in very general settings without structural assumptions on the unknown distribution, classical distances between probability measures might not the right choice. While creating an image that is different from the training images appears as creative and is favorable, the performance of the corresponding distribution in the Wasserstein $1$ distance is not necessarily good. In context of GANs, an inclusion of the distance between the generated distribution and the empirical measure has recently been investigated by Vardanyan et al. (2024).
@@ 5 Proofs
Before we begin with the proofs, let us introduce some additional notation: The spectral norm of $A$ is denoted by $\|A\|_{2}.$ We define the ceiling of $x\in\mathbb{R}$ as $\lceil x\rceil:=\min\{k\in\mathbb{Z}|k\geq x\}.$
For $\alpha\in(0,1]$, we define the Hölder space $\mathcal{H}^{\alpha}$ by
$\mathcal{H}^{\alpha}(\Omega)\coloneqq\big\{f\colon\Omega\rightarrow\mathbb{R}^{d}\colon\|f\|_{\mathcal{H}^{\alpha}(\Omega)}<\infty\big\},\quad\text{ where }\quad\|f\|_{\mathcal{H}^{\alpha}(\Omega)}\coloneqq\sup_{x,y\in\Omega}\frac{|f(x)-f(y)|}{|x-y|^{\alpha}}.$ (18)
@@
5.1 Proofs of Section˜2
Proof of Lemma˜2.1.
@@ 5.1 Proofs of Section˜2
$\mathbb{E}_{X_{t}\sim p^{n}_{t}}[\langle\hat{v}_{t}(X_{t}),v^{n}_{t}(X_{t})\rangle]$ $=\int\langle\hat{v}_{t}(x),v^{n}_{t}(x)\rangle p^{n}_{t}(x)\;\mathrm{d}x$
@@
5.2 Proofs and extended results of Section˜3
Proof of Example˜3.1.
For the Wasserstein $1$ distance, we can bound
$\mathsf{W}_{1}(\mathbb{P}^{\psi_{1}^{(1)}(Z)},\mathbb{P}^{\psi_{2}^{(1)}(Z)})$ $\leq\mathbb{E}_{Z\sim\mathcal{N}(0,1)}[|\psi_{1}^{(1)}(Z)-\psi_{1}^{(2)}(Z)|]$
$=\mathbb{E}_{Z\sim\mathcal{N}(0,1)}[|Z-\psi_{1}^{(2)}(Z)|]$
$=\mathbb{E}_{Z\sim\mathcal{N}(0,1)}\Big[\Big|\int_{0}^{1}v_{t}^{(2)}(\psi^{(2)}_{t}(Z))\;\mathrm{d}t\Big|\Big]$
$=\mathbb{E}_{Z\sim\mathcal{N}(0,1)}[|\varepsilon\sin\Big(\frac{\psi^{(2)}_{t}(Z)}{\varepsilon}\Big)|]$
$\leq\varepsilon.$
For the total variation distance, we first note that by the change of variables theorem, we have for every element of a sequence of diffeomorophisms $(f_{k})_{k\in\mathbb{N}}$
$f_{k}(Z)\sim\frac{\varphi(f_{k}^{-1}(\cdot))}{|f_{k}^{\prime}(f_{k}^{-1}(\cdot))|},$
where $\varphi$ is the density of $\mathcal{N}(0,1)$ and where we identified the density with the distribution. By definition of $v^{(1)}$, we have that $\psi^{(1)}_{t}(Z)\sim\varphi$. In case $f_{k}^{\prime}(x)>0$ for all $x\in\mathbb{R}$, we obtain for the total variation distance
$\operatorname{TV}(\mathcal{N}(0,1),\mathbb{P}^{f_{k}(Z)})$ $=\frac{1}{2}\int\Big|\varphi(x)-\frac{\varphi(f_{k}^{-1}(\cdot))}{|f^{\prime}(f_{k}^{-1}(\cdot))|}\Big|\;\mathrm{d}x$
$=\frac{1}{2}\int|\varphi(f_{k}(x))f_{k}^{\prime}(x)-\varphi(x)|\;\mathrm{d}x$
$=\frac{1}{2}\int|\varphi(x)(f_{k}^{\prime}(x)-1)+f_{k}^{\prime}(x)(\varphi(f_{k}(x))-\varphi(x))|\;\mathrm{d}x$
$\geq\frac{1}{2}\int|\varphi(x)(f_{k}^{\prime}(x)-1)|\;\mathrm{d}x-\frac{1}{2}\int|f_{k}^{\prime}(x)(\varphi(f_{k}(x))-\varphi(x))|\;\mathrm{d}x.$ (19)
In case $\sup_{k\in\mathbb{N}}\|f_{k}^{\prime}\|_{\infty}\leq C<\infty$ and $\|f_{k}-\operatorname{Id}\|_{\infty}\rightarrow 0$, we have that for $k$ large enough
$|f_{k}(x)-x|\leq 1\quad\Longrightarrow\quad|f_{k}(x)|\geq|x|-1,$
which implies
$|f_{k}^{\prime}(x)(\varphi(f_{k}(x))-\varphi(x))|\leq C\Big(\frac{1}{\sqrt{2\pi}}e^{-\frac{(|x|-1)^{2}}{2}}+\varphi(x)\Big).$
Since the right hand side of the above equation is integrable, we can use dominated convergence and the continuity of $\varphi$ to conclude that the second term in (19) converges towards 0 for $k\rightarrow\infty$. However, for the first term, we also need that $f^{\prime}\rightarrow(\operatorname{Id})^{\prime}=1$. In the Flow Matching setting, this corresponds to the derivative of $\psi^{(2)}_{1}$ with respect to $x$ converging to $1$, which is not controlled via the Flow Matching objective.
In the given example, we can solve the corresponding ODEs and obtain
$\frac{\partial\psi^{(1)}_{t}(x)}{\partial t}=0,\quad$ $\psi^{(1)}_{0}(x)=x,$ $\Longrightarrow\quad$ $\psi^{(1)}_{t}(x)=x$
$\frac{\partial\psi^{(2)}_{t}(x)}{\partial t}=\varepsilon\sin\Big(\frac{\psi^{(2)}_{t}(x)}{\varepsilon}\Big),\quad$ $\psi^{(2)}_{0}(x)=x,$ $\Longrightarrow\quad$ $\psi^{(2)}_{t}(x)=2\varepsilon\arctan\Big(e^{t}\tan\Big(\frac{x}{2\varepsilon}\Big)\Big)+2\pi\varepsilon\Big\lfloor\frac{x}{2\pi\varepsilon}+\frac{1}{2}\Big\rfloor,$
where $\psi^{(2)}_{t}$ is defined for almost all $x\in\mathbb{R}$. Note that $\big\lfloor\frac{x}{2\pi\varepsilon}+\frac{1}{2}\big\rfloor$ is chosen such that we obtain a continuous, differentiable and bijective function on $\mathbb{R}$. The solution of $\psi^{(2)}$ can be checked calculating the derivative with respect to $t$ and using $\sin(2\theta)=\frac{2\tan(\theta)}{1+\tan^{2}(\theta)}$. The derivative with respect to $x$ of $\psi_{t}^{(2)}$ is given by
$\frac{\partial\psi^{(2)}_{t}}{\partial x}=\frac{e^{t}\sec^{2}\big(\frac{x}{2\varepsilon}\big)}{e^{2t}\tan^{2}\big(\frac{x}{2\varepsilon}\big)+1}=\frac{e^{t}}{\big(e^{2t}-1\big)\sin^{2}\big(\frac{x}{2\varepsilon}\big)+1}>0.$
Since $t\in[0,1]$, the derivative is bounded. Therefore, we can use the lower bound from (19). By construction $\psi_{1}^{(2)}(x)\rightarrow x$ for $\varepsilon\rightarrow 0$, thus the second term in (19) converges towards 0. For the second term
$\int\Big|$ $\varphi(x)\Big(\frac{e}{\big(e^{2}-1\big)\sin^{2}\big(\frac{x}{2\varepsilon}\big)+1}-1\Big)\Big|\;\mathrm{d}x$
$\geq\frac{1}{\sqrt{2\pi e}}\int_{-1}^{1}\Big|\Big(\frac{e}{\big(e^{2}-1\big)\sin^{2}\big(\frac{x}{2\varepsilon}\big)+1}-1\Big)\Big|\;\mathrm{d}x$
$\geq\frac{1}{2\sqrt{2\pi e}}\int_{-1}^{1}\mathds{1}\Big\{\frac{e}{\big(e^{2}-1\big)\sin^{2}\big(\frac{x}{2\varepsilon}\big)+1}\leq\frac{1}{2}\Big\}\;\mathrm{d}x$
$=\frac{1}{2\sqrt{2\pi e}}\int_{-1}^{1}\mathds{1}\Big\{x\in\bigcup_{n\in\mathbb{Z}}\Big[2\varepsilon\Big(\pi n+\arcsin\Big(\sqrt{\frac{2e-1}{e^{2}-1}}\Big)\Big),2\varepsilon\Big(\pi(n+1)-\arcsin\Big(\sqrt{\frac{2e-1}{e^{2}-1}}\Big)\Big)\Big]\Big\}\;\mathrm{d}x.$
Denoting $\gamma\coloneqq\arcsin\Big(\sqrt{\frac{2e-1}{e^{2}-1}}\Big)$, the length of the intervals in the bound is $2\varepsilon((n+1)\pi-\gamma)-2\varepsilon(n\pi+\gamma)=2\varepsilon(\pi-2\gamma).$ Next we count how many intervals are lie completely in [1,1][-1,1]. We have that
$2\varepsilon(\pi n+\gamma)\geq-1\Longleftrightarrow n\geq-\frac{1}{2\varepsilon\pi}-\frac{\gamma}{\pi},\quad 2\varepsilon(\pi(n+1)-\gamma)\leq 1\Longleftrightarrow n\leq\frac{1}{2\varepsilon\pi}+\frac{\gamma}{\pi}-1.$
Therefore, we have
$\Big\lfloor\frac{1}{2\varepsilon\pi}+\frac{\gamma}{\pi}-1\Big\rfloor-\Big\lceil-\frac{1}{2\varepsilon\pi}-\frac{\gamma}{\pi}\Big\rceil$
intervals that lie completely in [1,1][-1,1]. We bound this number via
$\Big\lfloor\frac{1}{2\varepsilon\pi}+\frac{\gamma}{\pi}-1\Big\rfloor-\Big\lceil-\frac{1}{2\varepsilon\pi}-\frac{\gamma}{\pi}\Big\rceil\geq\frac{1}{2\varepsilon\pi}+\frac{\gamma}{\pi}-2-(l-\frac{1}{2\varepsilon\pi}-\frac{\gamma}{\pi}+1)=\frac{1}{\varepsilon\pi}+\frac{\gamma}{\pi}-3.$
Thus, we can bound
$\int\Big|\varphi(x)\Big(\frac{e}{\big(e^{2}-1\big)\sin^{2}\big(\frac{x}{2\varepsilon}\big)+1}-1\Big)\Big|\;\mathrm{d}x$ $\geq\frac{1}{2\sqrt{2\pi e}}2\varepsilon(\pi-2\gamma)\big(\frac{1}{\varepsilon\pi}+\frac{\gamma}{\pi}-3\big)$
$=\frac{1}{2\sqrt{2\pi e}}\Big(\frac{2}{\pi}(\pi-2\gamma)+2\varepsilon\gamma\frac{\pi-2\gamma}{\pi}-6\varepsilon(\pi-2\gamma)\Big)$
$\overset{\varepsilon\rightarrow 0}{\longrightarrow}\frac{\pi-2\gamma}{\pi\sqrt{2\pi e}}.$
Setting $0<c<\frac{\pi-2\gamma}{\pi\sqrt{2\pi e}}$ yields the result. ∎
Proof of Theorem˜3.2.
As $\mathsf{W}_{1}$ and $\mathsf{W}_{2}$ satisfy the triangle inequality (see Villani (2008, p. 94)),